# Strong Normalization for the Safe Fragment of a Minimal Rewrite Calculus: A Triple-Lexicographic Proof and a Conjecture on Full-Termination Limits for Pure Recursive Calculi (PRC)

### v4

@@ Abstract
Standard termination methods fail on this calculus due to a core structural barrier: the duplication of the step argument in the recursor. The rec-succ rule redistributes its argument into both the function application and the recursive call ($s\to f(s,\dots)$), an operation that algebraically defeats additive measures ($M_{\text{after}}\geq M_{\text{before}}$ whenever $M_{s}\geq 1$). Because of this duplication, thirteen distinct internal proof strategies succumb to the calculus’s complexity: additive counters, simple lexicographic measures, polynomial interpretations, single-bit flags, pure ordinal measures, structural depth counters, naive multiset orderings, local helper lemmas, complex hybrid measures, unchecked recursion, pure precedence orders, linear weight functions (KBO-style), and standard tree depth. Furthermore, classical path orders (LPO/MPO) bypass this barrier by relying on the universal Subterm Property, which acts as an imported external axiom. Counting these external path-order methods yields fourteen total strategy families (Section 7). This pattern indicates a structural limitation for calculi of this form rather than a defect of proof search.
@@
We state a central conjecture: For Pure Recursive Calculi (PRC), no internally definable method currently known to us proves full-calculus termination. Unlike Turing machines (which require an external tape), lambda calculus (which requires variable binding), and recursive functions (which import zero and successor as external primitives), a PRC has none of these external scaffolds. On this framing, the termination barrier is positioned at a more primitive layer than those classical settings.
These findings motivate a form of Pre-Arithmetic Incompleteness. Unlike Gödel’s theorems, which require arithmetic sufficient for encoding (Robinson Q), the barrier studied here appears at a more primitive layer: in operator calculi possessing only order and recursion, before arithmetic emerges. In this setting, the difficulty of proving full termination from internal methods parallels halting-style limitations at the substrate of operator definitions.
@@ 1 Introduction
This paper presents a minimal operator-only rewrite calculus (KO7) where the object language contains only constructors and operators with rewrite rules. There are no binders, types, external axioms, or semantic predicates; the rules are the semantics. The primary goals are: (i) a clean proof of strong normalization (SN) that handles duplication; (ii) a certified normalizer that always returns a normal form; (iii) unique normal forms via Newman’s Lemma, with the required local-confluence hypothesis discharged in the artifact; and (iv) a conjecture regarding the unprovability of termination for Pure Recursive Calculi (PRC) using internal methods.
@@
A second conceptual goal is to situate KO7 against results about fixed-target reachability in terminating calculi. If an internal provability predicate is defined by “$t$ reduces to $\top$”, then under SN the set decidability of $\{t\mid t\to^{*}\top\}$ is reducible to normalization. With confluence, decision reduces to normal-form equality.
@@
Strong normalization is proven for a guarded fragment of KO7 using a triple-lexicographic measure that combines a phase bit, multiset ordering (Dershowitz-Manna), and a computable natural rank. This result supports a certified normalizer with proven totality and soundness. Furthermore, a Newman engine for the safe fragment is formalized, and the global local-join hypothesis is fully discharged, yielding unconditional confluence and unique normal forms for the safe fragment.
Furthermore, we apply established proof theory to classify the boundaries of internal termination proving. Drawing on the work of Ferreira and Zantema [8], Buchholz [3], Gallier [9], and Rathjen–Weiermann [12], we observe that classical simplification orders (like LPO and MPO) fall outside our internal-definability boundary for pure calculi. These methods achieve termination by introducing the universal Subterm Property ($f(t)>t$) as a precondition, which in proof-theoretic analyses is tied to Kruskal-style well-foundedness arguments and transfinite induction strength (e.g., $\Pi^{1}_{1}$-comprehension). Because these orders import external structural axioms and well-ordering principles, they are outside the internal method class used in this paper. While not a proof of general non-termination, the work includes formally verified impossibility results for fourteen distinct strategy families under duplication. These machine-checked barriers provide evidence of structural limits in internal termination proving for self-referential calculi, directly motivating our central conjecture.
Highlights (formalization summary)
@@ 2 Background: Term Rewriting, SN, reachability, and Newman
Fixed-target (“small-term”) reachability in terminating calculi has well-charted complexity: NP-complete for length-reducing calculi (dropping to P under confluence), NExpTime/N2ExpTime for (linear) polynomial interpretations, and PSPACE for KBO-terminating calculi [1]. Modularity holds in certain linear, non-collapsing combinations [4], while even flat non-linear calculi exhibit undecidable reachability and confluence [10]. Termination under duplication typically requires orders beyond plain sizes; a naive size can increase under duplicating rules. The cure is a multiset extension of a base order [5] or (recursive) path orders [7].
@@ 3 The KO7 calculus
Table 1: KO7’s 8 kernel rules. In the full kernel relation, R_eq_refl and R_eq_diff overlap at eqW a a. The certified SafeStep fragment adds guards: R_eq_diff requires $a\neq b$, R_merge_void_left/right require $\delta$-flag tie conditions, R_merge_cancel requires both $\delta$-flag and $\kappa^{M}$ guards, R_rec_zero requires a $\delta$-flag guard on $b$, and R_eq_refl requires $\kappa^{M}(a)=0$. R_merge_cancel is collapsing. (* Note: R_rec_succ duplicates the step argument $s$ (one occurrence on the LHS, two on the RHS). This duplication is the termination stressor: the paper proves strong normalization for the guarded SafeStep fragment, not for the unrestricted full relation.)
@@
The full kernel relation Step is generated by the 8 rules in Table 1, while the certified artifact is proved for a guarded subrelation SafeStep.
@@
    • If $a$ and $b$ are syntactically identical ($a=b$), then $\mathrm{eqW}\,a\,b\Rightarrow\mathrm{void}$ (provided $\kappa^{M}(a)=0$).
@@ 4 Strong normalization
Genealogy of failures (why DM/MPO)
@@
  • Additive counters $\rho$: by the identity $\rho(\text{after})=\rho(\text{before})-1+\rho(S)$ there is no strict drop when $\rho(S)\geq 1$.
@@ 5 A certified normalizer
A normalization function is defined by well-founded recursion on SafeStepRev, with well-foundedness certified by the computable measure stack $\mu^{3}_{c}$. The normalizer operates on the root step relation SafeStep (not the context closure SafeStepCtx); a term is in normal form when no root safe rule applies ($\mathrm{NormalFormSafe}(t)\iff\neg\exists u.\;\mathrm{SafeStep}\;t\;u$). Properties proven in the formalization include:
@@ 6 Local confluence and Newman (guarded safe relation)
Scope and Guarantees (formalization-accurate)
The formal results concern a guarded safe subrelation SafeStep for which we prove:
  • SN (SafeStep). The computable KO7 triple measure $\mu^{3}_{c}$ strictly drops on every SafeStep. This yields a certified normalizer that is total and sound for the safe fragment.
  • Local-Join Library (SafeStep). Local-join lemmas per root shape and context wrappers are provided; the global hypothesis $\mathsf{locAll}$ is discharged in the artifact (locAll_safe), yielding unconditional confluence and unique NFs.
@@ 7 Impossibility results: The Fourteen Failed Methods
    Additive Counters: $\kappa(t)+k$ fails for any fixed $k$. The duplication rule $M(\text{after})=M(\text{before})-1+M(s)$ implies no strict drop when $M(s)\geq 1$.
@@
We distinguish between what is machine-checked in the code and what is argued in the meta-theory regarding Path Orders (Item 11). The Lean artifact (Conjecture_Boundary.lean) formally demonstrates that the isolated components of path orders (pure precedence, Item 12, and linear weights, Item 13) mathematically fail to orient the calculus. In classical proofs, successful path-order arguments rely on the universal Subterm Property ($f(t)>t$). While full LPO is a classical external method that could theoretically orient such systems, doing so imports this property. As argued in Section 8.2, this is a meta-theoretical intervention (via Kruskal-style assumptions) rather than an internally definable proof under Definition 9.1. The code demonstrates pressure toward such an axiom; the text analyzes its scope.
@@ 8 Decidability of Reachability
This connects the work to classical results: encoding undecidable properties through reduction to constants violates known theoretical limits. This motivates stratified approaches in proof assistants where object-level and meta-level reasoning are kept separate.
Assumptions (model)
The model assumes a finite first-order rewrite calculus over finite terms. The decision procedure computes a normal form (by SN) and checks whether it is $\top$; with confluence (hence unique normal forms), this reduces to one normalization and an equality test.
Complexity context
Small-term reachability in terminating calculi ranges from NP (length-reducing) to NExpTime/N2ExpTime (polynomial interpretations) and PSPACE (KBO), with confluence lowering the length-reducing class to P [1]. This situates our “normalize and compare” decision procedure for KO7 within the known complexity results.
Moreover, decidability can be modular for disjoint unions under left-linearity/non-collapsing assumptions [4], while termination alone does not guarantee decidability: even flat non-linear calculi have undecidable reachability/joinability/confluence [10].
@@ 9.1 The Sequential Nature of KO7 and Recursors
is not merely a reduction; it is a program. It applies the “step” term $s$ repeatedly, $n$ times. This is functionally identical to the recursor in Gödel’s System T or the step function of primitive recursive arithmetic. Both mechanisms allow a calculus to execute a sequence of operations of unbounded length.
@@
This duplication is the structural barrier. As demonstrated in the Impossibility Results (Section 7) and formally witnessed in the artifact by the theorem rec_succ_additive_barrier, this redistribution of $s$ defeats additive termination measures ($M_{\text{after}}\geq M_{\text{before}}$ whenever $M_{s}\geq 1$). A calculus that internalizes control flow in this way faces a sharp difficulty when proving full termination with simple internal metrics, because the metric must account for the unbounded potential of the duplicated step function.
@@
In this PRC setting, ordered computation is represented with a recursor. The recursor redistributes its step argument across recursive calls. This redistribution defeats additive termination measures: $M(\text{after})=M(\text{before})-1+M(s)$, yielding no strict drop when $M(s)\geq 1$.
Definition 9.1 (Internally definable measure).
An internally definable termination method for signature $\Sigma$ uses only: DM-multiset lifts of $\mathbb{N}$-valued ranks, and algebraic interpretations definable without importing external axioms. No encodings into external arithmetic, no borrowed logic, and no hard-coded constants outside the operator language.
Importantly, an internally definable method must be derived strictly from the operational semantics (the rewrite rules). It cannot inject universal structural axioms from the meta-theory. For example, standard simplification orders (LPO/RPO/MPO) are explicitly excluded because they rely on the universal Subterm Property ($f(t)>t$). As demonstrated by Ferreira and Zantema [8], this property acts as an imported precondition axiom, artificially establishing a fixed point (an anchor) that the pure calculus does not possess. Furthermore, as established in the proof-theoretic literature (e.g., Buchholz [3], Dershowitz & Okada [6]), the well-foundedness of these orders relies on Kruskal’s Tree Theorem, which Gallier [9] notes is a classic independence phenomenon requiring transfinite induction at or above $\varepsilon_{0}$. The distinction is: an internally definable method must be self-contained within the operator language; any proof that borrows structural axioms or well-ordering strength from an external theory is performing meta-level reasoning, not internal reasoning.
@@
KO7 with full Step is a PRC. Classical Turing machines fail condition 2 (they require an external tape). The untyped lambda calculus fails condition 1 ($\lambda$-abstraction introduces variable binding). Systems of primitive recursive functions fail condition 3 (zero and successor are imported as external primitives, not operator-defined). In the sense of this definition, the class targets a more primitive setting than these classical computational models: the termination barrier identified here arises before arithmetic, variable binding, or external memory enter the picture.
@@ 9.2 The Subterm Axiom and Artificial Fixed Points
9.3 The Pre-Arithmetic Incompleteness Hierarchy
The conjecture situates KO7 within a three-layer hierarchy of incompleteness results, presented as requiring progressively more external structure:
Table 3: Three layers of computational incompleteness. PRC (Layer 0) requires only an object-level recursor and no external scaffolding; Turing machines (Layer 1) add an external tape and meta-level iteration; Gödel systems (Layer 2) additionally require arithmetic encoding.
Layer Calculus class Minimum requirement Barrier
0 (this work) PRC Order + object-level recursor No known internally definable proof of full termination
1 (Turing) General computation External tape + meta-level transition Cannot decide halting (undecidable)
2 (Gödel) Arithmetic calculus Robinson Q (++ and $\times$) Cannot prove own consistency internally
Why Layer 0 Is Placed Below Layer 1
A classical Turing machine performs sequential computation via a meta-level transition function $\delta:Q\times\Gamma\to Q\times\Gamma\times\{L,R\}$. The iteration (“run for $k$ steps”) is defined outside the machine as a sequence of configuration yields. The tape is external memory; there is no operator within the term language that internalizes the step. In a PRC, by contrast, the recursor $\mathrm{rec}\Delta$ is an object-level constructor: “do the next step” is a rewrite rule applied from inside the term, not a meta-operation on a configuration. This is why a Turing machine fails condition 2 of Definition 9.2 and falls outside the PRC class.
Why Layer 0 Is Placed Below Layer 2
Gödel’s incompleteness theorems require arithmetic sufficient for encoding, at minimum Robinson arithmetic $Q$, with addition, multiplication, zero, and successor as imported primitives. Presburger arithmetic (addition only) and Skolem arithmetic (multiplication only) are both decidable and complete; they escape Gödel because they lack the encoding power of full $Q$. A PRC requires none of this: no imported arithmetic, no Gödel numbering, no self-referential encoding. The termination barrier at Layer 0 arises from structural duplication in the recursor (a constructive phenomenon) rather than from a diagonal self-reference argument. Arithmetic itself is a higher-layer phenomenon: it emerges when the PRC operators are later interpreted (e.g., $\delta$ as successor, $\mathrm{rec}$ as primitive recursion). The Layer 0 barrier exists before that interpretation is imposed. This explains why single-level Gödel encodings cannot coexist with globally terminating proof search; stratification is required.
@@ 9.3 The Pre-Arithmetic Incompleteness Hierarchy
  • Additive and polynomial methods are formally excluded in the audited family. The redistribution identity $M(\text{after})=M(\text{before})-1+M(s)$ yields no strict drop when $M(s)\geq 1$; polynomial interpretations without imported arithmetic degrade to additive measures and fail by the same identity (Lemma 4.2). Machine-checked barriers in Meta/Conjecture_Boundary.lean formally witness failure of additive counters, simple lexicographic pairs, flag-only methods, polynomial multiplicative ties, structural depth, and constellation-style hybrids (Section 7).
@@
    Note on orientation layers: The canonical SafeStep termination proof in this paper is the computable stack $(\delta,\kappa^{M},\tau)$ from Meta/ComputableMeasure.lean: rec_zero is discharged by DM on $\kappa^{M}$, rec_succ by the $\delta$ drop, and merge_cancel/eqW tie branches by $\tau$. MPO/LPO appear only as external meta-theoretic comparators in the boundary discussion, not as the internal certified certificate.
@@
In KO7, the rec-succ rule has the critical shape: the step argument $s$ appears twice on the RHS, and when $s$ contains nested redexes, additive measures fail to decrease strictly. The redistribution identity $M(\text{after})=M(\text{before})-1+M(s)$ yields no strict drop when $M(s)\geq 1$. The guarded SafeStep fragment terminates because the $\delta\text{-flag}$ restricts allowable step arguments; without such restriction, the full calculus admits unbounded recursive depth. At present, no internally definable ranking-function closure is known in this setting without external methods (axioms, encodings, or borrowed arithmetic).

### v5

@@ Abstract
Standard termination methods fail on this calculus due to a core structural barrier: duplication of the step argument in the recursor. The rec-succ rule redistributes its step argument into both the function application and the recursive call ($s\to f(s,\dots)$), which blocks strict decrease for additive families (formalized in Section 4). Because of this duplication, thirteen distinct internal proof strategies succumb to the calculus’s complexity: additive counters, simple lexicographic measures, polynomial interpretations, single-bit flags, pure ordinal measures, structural depth counters, naive multiset orderings, local helper lemmas, complex hybrid measures, unchecked recursion, pure precedence orders, linear weight functions (KBO-style), and standard tree depth. Furthermore, classical path orders (LPO/MPO) bypass this barrier by relying on the universal Subterm Property, which acts as an imported external axiom. Counting these external path-order methods yields fourteen total strategy families (Section 7). This pattern indicates a structural limitation for calculi of this form rather than a defect of proof search.
@@
We state a central conjecture (Section 9): For Pure Recursive Calculi, no internally definable method currently known to us proves full-calculus termination. The formally verified impossibility results and the conjecture boundary are detailed in Sections 7 and 9.
@@ 1 Introduction
This paper presents a minimal operator-only rewrite calculus (KO7) where the object language contains only constructors and operators with rewrite rules. There are no binders, types, external axioms, or semantic predicates; the rules are the semantics. The primary goals are: (i) a clean proof of strong normalization (SN) that handles duplication; (ii) a certified normalizer that always returns a normal form; (iii) unique normal forms via Newman’s Lemma, with the required local-confluence hypothesis discharged in the artifact; and (iv) a conjecture regarding the limits of internal termination proving for Pure Recursive Calculi (PRC).
@@
A second conceptual goal is to situate KO7 against results about fixed-target reachability in terminating calculi. If an internal provability predicate is defined by “$t$ reduces to a target constant cc”, then under SN the decidability of $\{t\mid t\to^{*}c\}$ is reducible to normalization. With confluence, decision reduces to normal-form equality.
@@
The main contributions, formally verified in the artifact, are:
@@
We additionally apply established proof theory to classify the boundaries of internal termination proving. Drawing on Ferreira and Zantema [8], Buchholz [3], Gallier [9], and Rathjen–Weiermann [12], we observe that classical simplification orders (like LPO and MPO) fall outside our internal-definability boundary for pure calculi: they achieve termination by importing the universal Subterm Property ($f(t)>t$), which in proof-theoretic analyses is tied to Kruskal-style well-foundedness arguments and transfinite induction strength (e.g., $\Pi^{1}_{1}$-comprehension). These machine-checked barriers provide evidence of structural limits in internal termination proving for self-referential calculi, directly motivating our central conjecture.
@@ 2 Background: Term Rewriting, SN, reachability, and Newman
Fixed-target (“small-term”) reachability in terminating calculi has well-charted complexity: NP-complete for length-reducing calculi (dropping to P under confluence), NExpTime/N2ExpTime for (linear) polynomial interpretations, and PSPACE for KBO-terminating calculi [1]. Modularity holds in certain linear, non-collapsing combinations [4], while even flat non-linear calculi exhibit undecidable reachability and confluence [10]. Termination under duplication typically requires orders beyond plain sizes; a naive size can increase under duplicating rules. In classical rewriting, common responses are a multiset extension of a base order [5] and (recursive) path orders [7]; under our internal-method boundary (Definition 9.1), only the former is admitted after computable projection to numeric payloads, while path orders are treated as external meta-theoretic methods.
@@ 3 The KO7 calculus
Table 1: KO7’s 8 kernel rules. In full Step, R_eq_refl and R_eq_diff overlap at eqW a a. The SafeStep fragment adds per-rule guards ($\delta$-flag, $\kappa^{M}$, and $a\neq b$ conditions) to resolve this overlap and ensure termination. R_rec_succ (*) duplicates the step argument $s$; this duplication is the central termination stressor.
@@
The full kernel relation Step is generated by the 8 rules in Table 1. The certified artifact is proved for a guarded subrelation SafeStep $\subseteq$ Step, which adds per-rule side conditions ($\delta$-flag, $\kappa^{M}$, and disequality guards) that restrict when each rule may fire.
@@
    • If $a$ and $b$ are syntactically identical ($a=b$), then $\mathrm{eqW}\,a\,b\Rightarrow\mathrm{void}$ (in the safe fragment, provided $\kappa^{M}(a)=0$; unconditional in full Step).
@@ 4 Strong normalization
Genealogy of failures (why DM and guarded lex components)
@@
  • Additive counters $\rho$: by the duplication stress identity above, there is no strict drop when $\rho(S)\geq 1$.
@@ 5 A certified normalizer
[…]
@@ 7 Impossibility results: The Fourteen Failed Methods
    Additive Counters: $\kappa(t)+k$ fails for any fixed $k$. The duplicating recursor branch blocks strict additive decrease (Section 4).
@@
We distinguish between what is machine-checked in the code and what is argued in the meta-theory regarding Path Orders (Item 11). The Lean artifact (Conjecture_Boundary.lean) formally demonstrates that the isolated components of path orders (pure precedence, Item 12, and linear weights, Item 13) mathematically fail to orient the calculus. In classical proofs, successful path-order arguments rely on the universal Subterm Property ($f(t)>t$). While full LPO is a classical external method that could theoretically orient such systems, doing so imports this property. As discussed in Section 9, this is a meta-theoretical intervention (via Kruskal-style assumptions) rather than an internally definable proof under Definition 9.1. The code demonstrates pressure toward such an axiom; the text analyzes its scope.
@@ 8 Decidability of Reachability
For SafeStep, this is immediate from Cor. 4.4 and Theorem 6.1: compute $\mathrm{normalizeSafe}(t)$ and compare it with the target constant. This section is included only to record that reduction from reachability to normalization is standard in the terminating+confluent setting. Complexity bounds for related classes are surveyed in [1, 4, 10].
@@ 9 Conjecture: Full Termination Limits in Pure Recursive Calculi (PRC)
Sections 47 contain the proved technical core of this paper. The present section states a scoped conjectural boundary motivated by those formal results.
@@ 9.1 The Sequential Nature of KO7 and Recursors
is not merely a reduction; it is a program. It applies the “step” term $s$ repeatedly, $n$ times. This is structurally analogous to the recursor in Gödel’s System T, though KO7’s recursor is untyped and unrestricted. Both mechanisms allow a calculus to execute a sequence of operations of unbounded length.
@@
This duplication is the structural barrier. As demonstrated in Section 7 and formally witnessed by rec_succ_additive_barrier, this redistribution defeats strict additive decrease, so simple internal metrics fail to orient the unrestricted branch.
@@
In this PRC setting, ordered computation is represented with a recursor. The recursor redistributes its step argument across recursive calls, yielding the duplication non-drop mechanism formalized in Section 4.
Definition 9.1 (Internally Definable Method (design choice; testable)).
For a finite signature $\Sigma$ and rewrite relation $R$, this paper calls a termination method internally definable iff all of the following checks are met:
  1. 1.
    Computable payload extraction: a computable map $\nu:\mathrm{Terms}(\Sigma)\to\mathbb{N}\times(\mathrm{Multiset}\;\mathbb{N}\times\mathbb{N})$ is provided.
  2. 2.
    Admitted target order: decrease is measured only in the computable lexicographic order
[…]
  3. 3.
    Per-rule certificate: for every root rule instance $t\to_{R}t^{\prime}$, one proves $\nu(t^{\prime})<_{\mathrm{Lex}}\nu(t)$.
  4. 4.
    No imported universal constructor axioms: no global assumptions of the form “for all constructors $f$, f(u)>uf(u)>u” are used unless derivable from $R$ itself.
This boundary is an explicit methodological design choice for the conjecture scope in this paper. It defines which proof families are counted as “internal” in the audited comparison class.
[…]
@@
KO7 with full Step is a PRC. Classical Turing machines fail condition 2 (they require an external tape). The untyped lambda calculus fails condition 1 ($\lambda$-abstraction introduces variable binding). Systems of primitive recursive functions fail condition 3 (zero and successor are imported as external primitives, not operator-defined). In this scoped definition, the class isolates operator-level recursion without imported arithmetic or external memory.
@@ 9.2 The Subterm Axiom and Artificial Fixed Points
9.3 Layered perspective (discussion only)
For context, we include a layered comparison with Turing-style and arithmetic incompleteness frameworks in Appendix A.2. This comparison is intentionally conjectural discussion, not a theorem-level reduction or embedding result.
@@ 9.3 Layered perspective (discussion only)
  • Additive and polynomial methods are formally excluded in the audited family. By the duplication non-drop mechanism from Section 4, additive counters do not strictly decrease on unrestricted duplicating recursors; polynomial interpretations without imported arithmetic collapse to the same obstruction. Machine-checked barriers in Meta/Conjecture_Boundary.lean formally witness failure of additive counters, simple lexicographic pairs, flag-only methods, polynomial multiplicative ties, structural depth, and constellation-style hybrids (Section 7).
@@
    Note: The canonical SafeStep certificate is the computable stack $(\delta,\kappa^{M},\tau)$ from Meta/ComputableMeasure.lean (Section 4); MPO/LPO appear only as external comparators in the boundary discussion.
@@
In KO7, the rec-succ rule has the critical shape: the step argument $s$ appears twice on the RHS, and when $s$ contains nested redexes, additive measures fail to decrease strictly (Section 4). The guarded SafeStep fragment terminates because the $\delta$-phase and $\kappa^{M}$ guards restrict this branch; without such restriction, no internally definable ranking-function closure is currently known to us without external methods (axioms, encodings, or borrowed arithmetic).